# Simulating Polynomial-Time Nondeterministic Turing Machines via Nondeterministic Turing Machines

### v20

@@ 1. Introduction
[…]
@@ 1.1. Our Contributions
$\mathcal{P}\neq\mathcal{NP}$ [^2] So, from this paper, we know that resolving the $\mathcal{P}\overset{?}{=}\mathcal{NP}$ problem does not necessarily require diagonalization techniques.
@@ 1.2. Background and Prior Work
where $L$ is a decision problem, and $\overline{L}$ is the complement of $L$ (i.e., $\overline{L}=\Sigma^{*}\setminus L$ with the assumption that the language $L$ is over the alphabet $\Sigma$). That is to say, ${\rm co}\mathcal{NP}$ is the complement of $\mathcal{NP}$. Note that the complement of a decision problem $L$ is defined as the decision problem whose answer is “yes” whenever the input is a “no” input of $L$, and vice versa. In other words, for instance, according to Wikipedia [A3], ${\rm co}\mathcal{NP}$ is the set of decision problems where there exists a polynomial $p(n)$ and a polynomial-time bounded Turing machine $M$ such that for every instance $x$, $x$ is a no-instance if and only if for some possible certificate $c$ of length bounded by $p(n)$, the Turing machine $M$ accepts the pair $(x,c)$. To the best of our knowledge, the complexity class ${\rm co}\mathcal{NP}$ was introduced for the first time by Meyer and Stockmeyer [MS72] with the name “$\prod_{1}^{P}$,” and Stockmeyer wrote a full paper on the polynomial hierarchy, which also uses the notation ${\rm co}\mathcal{NP}$; see e.g., [Sto77].
@@ 2. Preliminaries
Let $\mathbb{N}_{1}=\{1,2,3,\cdots\}$, where $+\infty\notin\mathbb{N}_{1}$, i.e., the set of all positive integers. We also denote by $\mathbb{N}$ the set of natural numbers,333 For the natural number system and $\infty$ (i.e., +$\infty$+ ∞ here), we refer the reader to the interesting textbook [Tao22], in which Tao provides a full and philosophical treatment of them. i.e.,
=$\mathbb{N}\overset{\rm def}{=}\mathbb{N}_{1}\cup\{0\}.$
@@
    $I$ is the set of input symbols; $I\subseteq T$.
@@
$M(w)=1,\quad\text{(i.e., polynomial-time nondeterministic Turing machine $M$ accepts $w$).}$.
@@ 3. All Polynomial-time Nondeterministic Turing Machines Are Enumerable
    The nondeterministic Turing machine itself is encoded by concatenating in any order the codes for each of the quintuples in its next-move function. Additional $1$’s may be prefixed to the string if desired. The result will be some string of 00’s and $1$’s, beginning with $1$, which we can interpret as an integer.
@@
$110^{k}$ 1
so that the tuple $(M,k)$ should be the concatenation of the binary string representing $M$ itself followed by the order $110^{k}$ 1. Now the tuple $(M,k)$ is encoded as a binary string, which can be explained as an integer.
@@
In the following context, we often use the notation $\langle M\rangle$ to denote the shortest binary string that represents the same $n^{k}+k$ time-bounded nondeterministic Turing machine $(M,k)$. We should further remark that the shortest binary string $\langle M\rangle$ that represents the same $n^{k}+k$ time-bounded nondeterministic Turing machine $(M,k)$ is by concatenating in any order the codes for each of the quintuples in its next-move function followed by the order $110^{k}$ 1 without additional $1$ prefixed to it, from which it is clear that $\langle M\rangle$ may not be unique (since the order of codes of the quintuples may be different), but the lengths of all different $\langle M\rangle$ are the same.
@@ 4. Simulation of All Polynomial-time Nondeterministic Turing Machines
    cells each, the blocks being separated by a single cell holding a marker ##\##, i.e., there are
@@
    cells, initialized to all 00’s. Tape $3$ is used as a counter to count up to
@@
    cells on tape $3$, initialized to all 00’s. Tape $3$ is used as a counter to count up to
@@
(4.1)4.1( 4.1 ) $1^{|w|-m}\langle M_{i}\rangle\in L_{s}\quad\text{if and and only if}\quad 1^{|w|-m}\langle M_{i}\rangle\in L(M_{w})=L(M_{i}).$
@@
(4.2)4.2( 4.2 ) $1^{|w|-m}\langle M_{i}\rangle\in L(M_{i})\quad\text{if and only if}\quad 1^{|w|-m}\langle M_{i}\rangle\notin\overline{L}(M_{i}).$
@@ 5. Proving That $L_{s}\in\mathcal{NP}$
for a fixed $i\in\mathbb{N}_{1}$, meaning that $U$ turns itself off when the counter of tape $3$ exceeds
@@
(5.2)5.2( 5.2 ) $L_{s}^{1}\subseteq L_{s}^{2}\subseteq\cdots\subseteq L_{s}^{i}\subseteq L_{s}^{i+1}\subseteq\cdots$
@@
(5.3)5.3( 5.3 ) $L_{s}^{i}\in{\rm NTIME}[n^{i}]\subseteq\mathcal{NP},\quad\text{for any fixed $i\in\mathbb{N}_{1}$}.$
@@ 6. Breaking the “Relativization Barrier”
$M^{A}(w)=0,\quad\text{(i.e., polynomial-time nondeterministic oracle Turing machine $M^{A}$ rejects $w$)}$
@@
$M^{A}(w)=1,\quad\text{(i.e., polynomial-time nondeterministic oracle Turing machine $M^{A}$ accepts $w$).}$.
@@
There is an oracle A for which $\text{There is an oracle $A$ for which }\mathcal{P}^{A}=\mathcal{NP}^{A}.$
@@
    $U^{A}$ decodes the tuple encoded by $x$. If $x$ is not the encoding of some ${\rm co}\mathcal{NP}^{A}$ machine $M^{A}_{j}$ with oracle $A$ for some $j$ (by Definition 6.4, equivalently, $x$ can not be decoded into some polynomial-time nondeterministic oracle Turing machine $M^{A}_{j}$ for some $j$, i.e., $x$ represents the trivial oracle Turing machine with an empty next-move function), then GOTO $6$, else determine $t$, the number of tape symbols used by $M^{A}_{j}$; $s$, its number of states; and $k$, its order.  [^7] We suppose that the order $110^{k}$ 1 of $M^{A}_{j}$ for some $k\in\mathbb{N}_{1}$ is also encoded into the binary string representing ${\rm co}\mathcal{NP}^{A}$ machine $M^{A}_{j}$, which is similar to Section 3. The third tape of $U^{A}$ can be used as “scratch” memory to calculate $t$.
@@
    cells each, the blocks being separated by a single cell holding a marker ##\##, i.e., there are
@@
    cells, initialized to all 00’s. Tape $3$ is used as a counter to count up to
@@
    cells on tape $3$, initialized to all 00’s. Tape $3$ is used as a counter to count up to
@@
(6.1)6.1( 6.1 ) $1^{|w|-m}\langle M^{A}_{i}\rangle\in L^{A}_{s}\quad\text{if and only if}\quad 1^{|w|-m}\langle M^{A}_{i}\rangle\in L(M^{A}_{w})=L(M^{A}_{i}).$
@@
(6.2)6.2( 6.2 ) $1^{|w|-m}\langle M^{A}_{i}\rangle\in L(M^{A}_{i})\quad\text{if and only if}\quad 1^{|w|-m}\langle M^{A}_{i}\rangle\notin\overline{L}(M^{A}_{i}).$
@@
(6.5)6.5( 6.5 ) $L^{A,1}_{s}\subseteq L^{A,2}_{s}\subseteq\cdots\subseteq L^{A,i}_{s}\subseteq L^{A,i+1}_{s}\subseteq\cdots$
@@ Appendix A Two Definitions of $\mathcal{NP}$ Are Equivalent
$\begin{array}[]{ll}\mathcal{P}\overset{{\rm def}}{=}&\{L\,:\,L=L(M)\text{ for some Turing machine $M$ that runs}\\ &\qquad\text{ in polynomial time }\}.\end{array}$
@@
A.1. A Proof of Two Definitions of $\mathcal{NP}$ Are Equivalent.
@@
$\begin{array}[]{ll}\mathcal{NP}\overset{{\rm def}}{=}&\{L\,:\,L=L(M)\text{ forsome nondeterministic Turing machine $M$}\\ &\qquad\text{that runs in polynomial time}\}.\end{array}$
@@
$L\in\{\mathcal{L}\,:\,\mathcal{L}=L(M)\text{ for some nondeterministic Turing machine $M$ that runs in polynomial time}\}.$
@@
$L\in\{\mathcal{L}\,:\,\mathcal{L}=L(M)\,\text{ for some nondeterministic Turing machine $M$ that runs in polynomial time}\}.$
@@
It is time for us to show that Definition A.3 is equivalent to Definition A.1. For the purposes of brevity, we omit the proof and refer the reader to [Kar72] (cf. Theorem 1 in [Kar72]), or to the proof of Theorem $7$.20202020 in [Sip13], which is a theorem answering the same question discussed here (see p. $294$ of [Sip13]). Thus, this completes the proof.

### v21

@@ 1. Introduction
$\mathcal{NP}\overset{\rm def}{=}$ $\big{\{}L\,:\,\text{there exists some polynomial-time nondeterministic Turing machine}$
@@ 1.1. Our Contributions
$\mathcal{P}\neq\mathcal{NP}$. [^2] So, from this paper, we know that resolving the $\mathcal{P}\overset{?}{=}\mathcal{NP}$ problem does not necessarily require diagonalization techniques.
@@ 1.2. Background and Prior Work
where $L$ is a decision problem, and $\overline{L}$ is the complement of $L$ (i.e., $\overline{L}=\Sigma^{*}\setminus L$ with the assumption that the language $L$ is over the alphabet $\Sigma$). That is to say, ${\rm co}\mathcal{NP}$ is the complement of $\mathcal{NP}$. Note that the complement of a decision problem $L$ is defined as the decision problem whose answer is “yes” whenever the input is a “no” input of $L$, and vice versa. In other words, for instance, according to Wikipedia [A3], ${\rm co}\mathcal{NP}$ is the set of decision problems where there exists a polynomial $p(n)$ and a polynomial-time bounded Turing machine $M$ such that for every instance $x$, $x$ is a no-instance if and only if for some possible certificate $c$ of length bounded by $p(n)$, the Turing machine $M$ accepts the pair $(x,c)$. To the best of our knowledge, the complexity class ${\rm co}\mathcal{NP}$ was introduced for the first time by Meyer and Stockmeyer [MS72] with the name $\prod_{1}^{P}$,” and Stockmeyer wrote a full paper on the polynomial hierarchy, which also uses the notation ${\rm co}\mathcal{NP}$; see e.g., [Sto77].
@@ 2. Preliminaries
Let $\mathbb{N}_{1}=\{1,2,3,\cdots\}$, where $\infty\notin\mathbb{N}_{1}$, i.e., the set of all positive integers. We also denote by $\mathbb{N}$ the set of natural numbers,333 For the natural number system and $\infty$ (i.e., $\infty$ here), we refer the reader to the interesting textbook [Tao22], in which Tao provides a full and philosophical treatment of them. i.e.,
$\mathbb{N}\overset{\rm def}{=}\mathbb{N}_{1}\cup\{0\}.$
@@
    II is the set of input symbols; $I\subseteq T$.
@@
$M(w)=1,\quad\text{(i.e., polynomial-time nondeterministic Turing machine $M$ accepts $w$).}$
@@ 3. All Polynomial-time Nondeterministic Turing Machines Are Enumerable
    The nondeterministic Turing machine itself is encoded by concatenating in any order the codes for each of the quintuples in its next-move function. Additional $1$’s may be prefixed to the string if desired. The result will be some string of 0’s and $1$’s, beginning with $1$, which we can interpret as an integer.
@@
$10^{k}1$
so that the tuple $(M,k)$ should be the concatenation of the binary string representing $M$ itself followed by the order $10^{k}1$. Now the tuple $(M,k)$ is encoded as a binary string, which can be explained as an integer.
@@
In the following context, we often use the notation $\langle M\rangle$ to denote the shortest binary string that represents the same $n^{k}+k$ time-bounded nondeterministic Turing machine $(M,k)$. We should further remark that the shortest binary string $\langle M\rangle$ that represents the same $n^{k}+k$ time-bounded nondeterministic Turing machine $(M,k)$ is by concatenating in any order the codes for each of the quintuples in its next-move function followed by the order $10^{k}1$ without additional $1$ prefixed to it, from which it is clear that $\langle M\rangle$ may not be unique (since the order of codes of the quintuples may be different), but the lengths of all different $\langle M\rangle$ are the same.
@@ 4. Simulation of All Polynomial-time Nondeterministic Turing Machines
    cells each, the blocks being separated by a single cell holding a marker #\#, i.e., there are
@@
    cells, initialized to all 0’s. Tape $3$ is used as a counter to count up to
@@
    cells on tape $3$, initialized to all 0’s. Tape $3$ is used as a counter to count up to
@@
(4.1) $1^{|w|-m}\langle M_{i}\rangle\in L_{s}\quad\text{if and and only if}\quad 1^{|w|-m}\langle M_{i}\rangle\in L(M_{w})=L(M_{i}).$
@@
(4.2) $1^{|w|-m}\langle M_{i}\rangle\in L(M_{i})\quad\text{if and only if}\quad 1^{|w|-m}\langle M_{i}\rangle\notin\overline{L}(M_{i}).$
@@ 5. Proving That $L_{s}\in\mathcal{NP}$
for a fixed $i\in\mathbb{N}_{1}$ (In fact, there is another version of the construction of $L_{s}^{{}^{\prime}i}$ (which satisfies $L_{s}^{{}^{\prime}i}\subseteq L_{s}^{i}$) for a fixed $i\in\mathbb{N}_{1}$ without adding a new tape to $U$ as a counter to count up to $n^{i+1}$, which is put into the Appendix B), meaning that $U$ turns itself off when the counter of tape $3$ exceeds
@@
(5.2) $L_{s}^{1}\subseteq L_{s}^{2}\subseteq\cdots\subseteq L_{s}^{i}\subseteq L_{s}^{i+1}\subseteq\cdots$
@@
(5.3) $L_{s}^{i}\in{\rm NTIME}[n^{i}]\subseteq\mathcal{NP},\quad\text{for any fixed $i\in\mathbb{N}_{1}$}.$
@@ 6. Breaking the “Relativization Barrier”
[…]
@@
[…]
@@
$\text{There is an oracle $A$ for which }\mathcal{P}^{A}=\mathcal{NP}^{A}.$
@@
    $U^{A}$ decodes the tuple encoded by $x$. If $x$ is not the encoding of some ${\rm co}\mathcal{NP}^{A}$ machine $M^{A}_{j}$ with oracle $A$ for some $j$ (by Definition 6.4, equivalently, $x$ can not be decoded into some polynomial-time nondeterministic oracle Turing machine $M^{A}_{j}$ for some $j$, i.e., $x$ represents the trivial oracle Turing machine with an empty next-move function), then GOTO $6$, else determine $t$, the number of tape symbols used by $M^{A}_{j}$; $s$, its number of states; and $k$, its order.  [^7] We suppose that the order $10^{k}1$ of $M^{A}_{j}$ for some $k\in\mathbb{N}_{1}$ is also encoded into the binary string representing ${\rm co}\mathcal{NP}^{A}$ machine $M^{A}_{j}$, which is similar to Section 3. The third tape of $U^{A}$ can be used as “scratch” memory to calculate $t$.
@@
    cells each, the blocks being separated by a single cell holding a marker #\#, i.e., there are
@@
    cells, initialized to all 0’s. Tape $3$ is used as a counter to count up to
@@
    cells on tape $3$, initialized to all 0’s. Tape $3$ is used as a counter to count up to
@@
(6.1) $1^{|w|-m}\langle M^{A}_{i}\rangle\in L^{A}_{s}\quad\text{if and only if}\quad 1^{|w|-m}\langle M^{A}_{i}\rangle\in L(M^{A}_{w})=L(M^{A}_{i}).$
@@
(6.2) $1^{|w|-m}\langle M^{A}_{i}\rangle\in L(M^{A}_{i})\quad\text{if and only if}\quad 1^{|w|-m}\langle M^{A}_{i}\rangle\notin\overline{L}(M^{A}_{i}).$
@@
(6.5) $L^{A,1}_{s}\subseteq L^{A,2}_{s}\subseteq\cdots\subseteq L^{A,i}_{s}\subseteq L^{A,i+1}_{s}\subseteq\cdots$
@@ Appendix A Two Definitions of $\mathcal{NP}$ Are Equivalent
[…]
@@
A.1. A Proof of Two Definitions of $\mathcal{NP}$ Are Equivalent
@@
[…]
@@
[…]
@@
[…]
@@
It is time for us to show that Definition A.3 is equivalent to Definition A.1. For the purposes of brevity, we omit the proof and refer the reader to [Kar72] (cf. Theorem 1 in [Kar72]), or to the proof of Theorem 77.2020 in [Sip13], which is a theorem answering the same question discussed here (see p. $294$ of [Sip13]). Thus, this completes the proof.
Appendix B A New Construction of $L_{s}^{{}^{\prime}i}$
Note that the language $L_{s}^{{}^{\prime}i}$ is accepted by $U$ also running at most $n^{i+1}$ steps (which means $L_{s}^{{}^{\prime}i}\in\mathcal{NP}$); this means that $U$ simulates those $n^{k}+k$ time-bounded nondeterministic Turing machines with order $k$ that $k\leq i$, since the simulation takes
$(n^{k}+k)\log(n^{k}+k)<n^{k+1}$
steps, which is required to be less than or equal to $n^{i+1}$ steps (all of such $k$ obviously satisfy that $k\leq i$, since it requires that $k+1\leq i+1$). Thus, for a fixed $i\in\mathbb{N}_{1}$, the language $L_{s}^{{}^{\prime}i}$ can be constructed as follows:
Since all ${\rm co}\mathcal{NP}$ machines are enumerable, let $U$ be a four-tape nondeterministic Turing machine that operates as follows on an input string $x$ of length $n$.
  1. (1)
  2. (2)
    If $k+1\leq i+1$, then GOTO $3$, else reject the input $x$.
  3. (3)
    Then $U$ lays off on its second tape $n$ blocks of
    $\lceil\log t\rceil$
    cells each, the blocks being separated by a single cell holding a marker #\#, i.e., there are
    $(1+\lceil\log t\rceil)n$
  4. (4)
    On tape $3$, $U$ sets up a block of
    $\lceil(k+1)\log n\rceil$
    cells, initialized to all 0’s. Tape $3$ is used as a counter to count up to
    $n^{k+1}.$
  5. (5)
    $n^{k+1}$
  6. (6)
    $\lceil 2\times\log n\rceil$
    cells on tape $3$, initialized to all 0’s. Tape $3$ is used as a counter to count up to
    $n^{2}.$
    $n^{2}$
    steps, so it is within $O(n)$. Note that the number of $2$ in
    $\lceil 2\times\log n\rceil$
    is fixed, i.e., it is a default setting (or it is by default).
Compared with the proof of Theorem 4.1, the construction of $U$ is only different in the red item 2.
It is easy to see that $L_{s}^{{}^{\prime}i}\subseteq L_{s}^{i}$ for some $i$, since for any $w\in L_{s}^{{}^{\prime}i}$ we have that $w\in L_{s}^{i}$, but for some $\in L_{s}^{i}, ww$ may not be in $L_{s}^{{}^{\prime}i}$ (we leave it to the reader to determine why). But, since all polynomial-time nondeterministic Turing machines will be simulated by $U$ when $i$ runs through all numbers in $\mathbb{N}_{1}$, we certainly have that
$L_{s}=\bigcup_{i\in\mathbb{N}_{1}}L_{s}^{{}^{\prime}i}.$